OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

(x + a)(x − a) = x^2 - a^2 and (x + a)(x + a) = x^2 + a^2 + 2ax